OpenStudy (anonymous):

Reduce to lowest terms. 3m^2-3n^2/6m-6n

OpenStudy (anonymous):

\[\frac{3m^{2}-3n^{2}}{6m-6n} = \frac{3(m^{2}-n^{2})}{6(m-n)} = \frac{3(m-n)(m+n)}{6(m-n)}\] After canceling out the m - n terms, you end up with: \[\frac{3}{6}(m+n) = \frac{m+n}{2}\]